# 00footnotetext: Received xxxx June xxxx

### v1

@@ Abstract
Using 20.3 $\mathrm{fb}^{-1}$ of $e^{+}e^{-}$ annihilation data collected at a center-of-mass energy of 3.773 $\,\mathrm{GeV}$ with the BESIII detector, we report an improved search for the radiative leptonic decay $D^{+}\to\gamma e^{+}\nu_{e}$. An upper limit on its partial branching fraction for photon energies $E_{\gamma}>10\leavevmode\nobreak\ \,\mathrm{MeV}$ is determined to be $1.2\times 10^{-5}$ at 90% confidence level, which excludes most current theoretical predictions. A sophisticated deep learning approach with thorough validation, based on the Transformer architecture, is implemented to efficiently distinguish the signal from massive backgrounds.
@@ 1 Introduction
The purely leptonic decays of charmed meson, $D^{+}\to l^{+}\nu_{l}\ (l=e,\ \mu,\ \tau)$, offer an ideal laboratory for testing the Standard Model. These decays offer one of the simplest and most well-understood probes of the $c\to d$ quark flavour-changing transition [1, 2, 3, 4]. However, these decays are subject to helicity suppression, with decay rates proportional to the square of the lepton mass. Consequently, the $D^{+}\to e^{+}\nu_{e}$ decay rate is highly suppressed compared to its muonic [5, 6] and tauonic [5, 7] counterparts, with a branching fraction (BF) expected to be around $10^{-8}$ and beyond the current experimental sensitivity [8, 9].
@@
×$\times\left|V_{cd}\right|^{2}E_{\gamma}^{3}m_{D^{+}}\left(1-\frac{2E_{\gamma}}{m_{D^{+}}}\right),$
where $\alpha_{em}$ is the fine-structure constant, $G_{F}$ is the Fermi coupling constant, $\left|V_{cd}\right|$ is the Cabibbo-Kobayashi-Maskawa matrix element [11, 12], $m_{D^{+}}$ is the nominal mass of the $D^{+}$ meson [13], and $F_{V}$ and $F_{A}$ are the form factors characterizing the non-perturbative quantum chromodynamics (QCD) dynamics. Various theoretical models have been proposed to investigate the $D^{+}\to\gamma e^{+}\nu_{e}$ decay, including quark models [14, 15, 16], perturbative QCD [17] and QCD factorization approaches [18, 19, 10]. Most theoretical predictions on its BF, $\mathcal{B}(D^{+}\to\gamma e^{+}\nu_{e})$, are at the order of $10^{-5}$, as summarized in Table 1.
@@
[…]
The only experimental result of $D^{+}\to\gamma e^{+}\nu_{e}$ comes from the BESIII experiment [20], based on 2.9 $\rm fb^{-1}$ of $e^{+}e^{-}$ annihilation data taken at the center-of-mass energy $\sqrt{s}=3.773\leavevmode\nobreak\ \,\mathrm{GeV}$. The upper limit on the partial BF of $D^{+}\to\gamma e^{+}\nu_{e}$ with $E_{\gamma}>10$ MeV was determined to be $3.0\times 10^{-5}$ at 90% confidence level (C.L.), approaching the range of theoretical predictions [16, 17, 18, 19, 10]. The larger dataset of 20.3 $\rm fb^{-1}$ of $e^{+}e^{-}$ annihilation recently collected at BESIII [21, 22, 23], which includes the previous 2.9  $\rm fb^{-1}$, provides an excellent opportunity to further constrain these theoretical models.
@@
An unprecedentedly powerful signal identification tool can be accomplished via a deep learning approach, which utilizes a deep neural network (DNN) to distinguish signal from backgrounds. Due to the flexible data representation and modern algorithms, this method exhibits powerful capabilities in recognizing and interpreting underlying relationships and hidden patterns within the topological structure of $e^{+}e^{-}$ annihilation events.
In this work, we present an improved search for $D^{+}\to\gamma e^{+}\nu_{e}$ by analyzing the 20.3 $\mathrm{fb}^{-1}$ data, enhancing signal sensitivity through a deep-learning-based signal identification method. Throughout this paper, charge conjugation is implied by default.
@@ 2 BESIII detector and Monte Carlo simulation
The BESIII detector [24] records symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring [25] in the center-of-mass energy range from 1.84 to 4.95 GeV, with a peak luminosity of $1.1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $\sqrt{s}=3.773\;\text{GeV}$. BESIII has collected large data samples in this energy region [26, 27, 28]. The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identification modules interleaved with steel. The charged-particle momentum resolution at $1\leavevmode\nobreak\ {\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the plastic scintillator TOF barrel region is 68 ps, while that in the end cap region was 110 ps. The end cap TOF system was upgraded in 2015 using multigap resistive plate chamber technology, providing a time resolution of 60 ps, which benefits 85% of the data used in this analysis [29, 30, 31].
Monte Carlo (MC) simulated data samples produced with a geant4-based [32] software package, which includes the geometric description of the BESIII detector and the detector response, are used to determine detection efficiencies and to estimate backgrounds. A signal MC sample of the decay $D^{+}\to\gamma e^{+}\nu_{e}$ is simulated according to Eq. (1), where the form factors are taken from Ref. [19]. The minimum energy of the radiative photon is set at 10 MeV to avoid the infrared divergence for soft photons. The simulation models the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc [33, 34]. The inclusive MC sample includes the production of $D\bar{D}$ pairs (including quantum coherence for the neutral $D$ channels), the non-$D\bar{D}$ decays of the $\psi(3770)$, the ISR production of the $J/\psi$ and $\psi(3686)$ states, and the continuum processes incorporated in kkmc [33, 34]. All particle decays are modelled with evtgen [35, 36] using branching fractions either taken from the Particle Data Group [13], when available, or otherwise estimated with lundcharm [37, 38]. Final state radiation from charged final state particles is incorporated using the photos package [39].
@@ 3 Initial data analysis
Taking advantage of the $D^{+}D^{-}$ pair production in $e^{+}e^{-}$ annihilation at $\sqrt{s}=3.773$ GeV, we employ a double-tag (DT) analysis method [40] to constrain the kinematics of $D^{+}\to\gamma e^{+}\nu_{e}$ and suppress non-$D^{+}D^{-}$ backgrounds. The $D^{-}$ meson is firstly reconstructed through six specific hadronic decay modes, $K^{+}\pi^{-}\pi^{-}$, $K^{+}\pi^{-}\pi^{-}\pi^{0}$, $K^{0}_{S}\pi^{-}$, $K^{0}_{S}\pi^{-}\pi^{0}$, $K^{0}_{S}\pi^{+}\pi^{-}\pi^{-}$ and $K^{+}K^{-}\pi^{-}$. Here, $K^{0}_{S}$ is reconstructed with a $\pi^{+}\pi^{-}$ pair and $\pi^{0}$ is reconstructed from two photons. These decay modes are referred to as tag modes, and the selected events are called single-tag (ST) events. In the presence of ST $D^{-}$ meson, the signal decay $D^{+}\to\gamma e^{+}\nu_{e}$ is then searched for in its recoil system, with selected events referred to as DT events. The BF of the signal decay is determined by
@@
where $N_{\rm DT}$ is sum of DT yields for all tag modes, and $N^{i}_{\rm ST}$, $\epsilon^{i}_{\rm ST}$, and $\epsilon^{i}_{\rm DT}$ are the ST yield, ST efficiency, and DT efficiency for the tag mode $i$, respectively.
[…]
The $K_{S}^{0}$ candidates are reconstructed from two oppositely charged tracks satisfying $V_{z}<$ 20 cm. The two charged tracks are assigned as $\pi^{+}\pi^{-}$ without imposing further PID criteria. They are constrained to originate from a common vertex and are required to have an invariant mass in the range of (487, 511) $\,\mathrm{MeV}/c^{2}$ corresponding to approximately ±$\pm 3\sigma$ of the detection resolution. The decay length of the $K^{0}_{S}$ candidate is required to be greater than twice the vertex resolution away from the IP.
Photon candidates are identified using showers in the EMC. The deposited energy of each shower must be more than 25 MeV in the barrel region ($|\cos\theta|<0.80$) and more than 50 MeV in the end cap region ($0.920.86<|\cos\theta|<$ 0.92). To exclude showers that originate from charged tracks, the angle subtended by the EMC shower and the position of the closest charged track at the EMC must be greater than 10 degrees as measured from the IP. To suppress electronic noise and showers unrelated to the event, the difference between the EMC time and the event start time is required to be within [0, 700] ns. The $\pi^{0}$ mesons are reconstructed using a kinematic fit from all combinations of two photon candidates by constraining their invariant mass to the nominal $\pi^{0}$ mass [13]. The invariant mass of two photons before the kinematic fit is required to be in the range of (115, 150) $\,\mathrm{MeV}/c^{2}$. The combinations with both photons from end cap EMC regions are rejected due to worse resolution.
After selecting the component particles, we reconstruct ST $D^{-}$ meson using their combinations. Two kinematic variables, the energy difference $\Delta E$ and the beam-constrained mass $M_{\rm BC}$, are calculated for signal extraction
@@
where $E_{\rm beam}$ is the beam energy, $\vec{p}_{D^{-}}$ and $E_{D^{-}}$ are the momentum vector and energy of $D^{-}$ candidate in the $e^{+}e^{-}$ rest frame, respectively. If multiple combinations are present in an event, the combination with the smallest $|\Delta E|$ is retained for further analysis for each tag mode and charge. To suppress combinatorial backgrounds, mode-dependent requirements on the $\Delta E$ are imposed on the accepted candidates within approximate ±$\pm 3.5\sigma$ range of the detection resolution, as listed in Table 3.
@@
The $\Delta E$ requirements, ST yields in data, ST and DT efficiencies for each tag mode in data. Note the efficiencies do not include the BFs of $\pi^{0}$ and $K^{0}_{S}$ subsequent decays. Tag mode $\Delta E$ (MeV) $N_{\rm{ST}}^{i}\ (\times 10^{3})$ $\epsilon_{\rm{ST}}^{i}\ (\%)$ $\epsilon_{\rm{DT}}^{i}\ (\%)$ $K^{+}\pi^{-}\pi^{-}$ (25, 24)2524(-25,\ 24)( - 25 , 24 ) $5552.8\pm 2.5$ $51.10$ 10.78 $K^{0}_{S}\pi^{-}$ (25, 26)2526(-25,\ 26)( - 25 , 26 ) $656.5\pm 0.8$ $51.42$ 10.69 $K^{+}\pi^{-}\pi^{-}\pi^{0}$ (57, 46)5746(-57,\ 46)( - 57 , 46 ) $1723.7\pm 1.8$ $24.40$ 5.42 $K^{0}_{S}\pi^{-}\pi^{0}$ (62, 49)6249(-62,\ 49)( - 62 , 49 ) $1442.4\pm 1.5$ $26.45$ 5.57 $K^{0}_{S}\pi^{-}\pi^{-}\pi^{+}$ (28, 27)2827(-28,\ 27)( - 28 , 27 ) $790.2\pm 1.1$ $29.59$ 6.62 $K^{+}K^{-}\pi^{-}$ (24, 23)2423(-24,\ 23)( - 24 , 23 ) $481.4\pm 0.9$ $40.91$ 8.96
[…]
@@
The signal candidates for $D^{+}\to\gamma e^{+}\nu_{e}$ are selected from the remaining charged tracks and showers, excluding those associated with the ST $D^{-}$ meson. We require only one track left with charge opposite to that of the ST $D^{-}$ meson, following the same selection criteria as those in the tag side. This track must be identified as a positron combining the PID information from $\mathrm{d}E/\mathrm{d}x$, TOF and EMC, satisfying $\mathcal{L}(e)>0.001$ and $\mathcal{L}(e)/[\mathcal{L}(e)+\mathcal{L}(\pi)+\mathcal{L}(K)]>0.8$ for the combined confidence levels. We also require at least one photon candidate satisfying the same selection criteria as used in the tag side. If multiple photon candidates pass the selection, the most energetic one is chosen as the signal radiative photon. To improve the kinematic resolution of the positron, final-state-radiation recovery is performed by adding the momenta of any remaining photon candidates back to the reconstructed momentum of positron candidate if their opening angles with respect to the IP are less than $5^{\circ}$.
To derive the information of the missing neutrino, a kinematic observable $U_{\rm miss}$ is defined as
@@
We also study $D^{+}\to\pi^{0}e^{+}\nu_{e}$ as a validation channel in this work. In reconstructing the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ events, additional DT selection criteria are applied after those mentioned above. At least one $\pi^{0}$ candidate is required, in which the one with minimum $\chi^{2}$ of the kinematic fit is chosen as the signal $\pi^{0}$ candidate. Another kinematic observable, $U^{\prime}_{\rm miss}$, is defined as
where $E^{\prime}_{\rm miss}=E_{\rm beam}-E_{e^{+}}-E_{\pi^{0}}$ and $\vec{p^{\prime}}_{\rm miss}=\vec{p}_{D^{+}}-\vec{p}_{e^{+}}-\vec{p}_{\pi^{0}}$ are calculated using the energy and momentum of the signal $\pi^{0}$ candidate, respectively.
Figure 3(a) presents the $U_{\rm miss}$ distributions for both data and inclusive MC samples after applying the above selections. The major backgrounds are from $D^{+}\to\pi^{0}e^{+}\nu_{e}$, $D^{+}\to K^{0}_{L}e^{+}\nu_{e}$, and $D^{+}\to K^{0}_{S}(\to\pi^{0}\pi^{0})e^{+}\nu_{e}$. Notably, the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background exhibits a non-trivial shape within the signal region. The high background level makes it difficult to re-

### v2

# 00footnotetext: Received 20 March 2025; Accepted 17 April 2025; Published online 18 April 2025

@@ Abstract
Using 20.3 $\mathrm{fb}^{-1}$ of $e^{+}e^{-}$ annihilation data collected at a center-of-mass energy of 3.773 $\mathrm{GeV}$ with the BESIII detector, we report on an improved search for the radiative leptonic decay $D^{+}\to\gamma e^{+}\nu_{e}$. An upper limit on its partial branching fraction for photon energies $E_{\gamma}>10\penalty 10000\ \,\mathrm{MeV}$ was determined to be $1.2\times 10^{-5}$ at a 90% confidence level; this excludes most current theoretical predictions. A sophisticated deep learning approach, which includes thorough validation and is based on the Transformer architecture, was implemented to efficiently distinguish the signal from massive backgrounds.
@@ 1 Introduction
The purely leptonic decays of charmed mesons, $D^{+}\to l^{+}\nu_{l}\ (l=e,\ \mu,\ \tau)$, offer an ideal laboratory for testing the Standard Model. These decays also offer one of the simplest and best understood probes of the $c\to d$ quark flavour-changing transition [1, 2, 3, 4]. However, these decays are subject to helicity suppression, with decay rates proportional to the square of the lepton mass. Consequently, the $D^{+}\to e^{+}\nu_{e}$ decay rate is highly suppressed compared to its muonic [5, 6] and tauonic [5, 7] counterparts, with a branching fraction (BF) expected to be approximately $10^{-8}$, beyond the current experimental sensitivity [8, 9].
@@
$\times\left|V_{cd}\right|^{2}E_{\gamma}^{3}m_{D^{+}}\left(1-\frac{2E_{\gamma}}{m_{D^{+}}}\right),$
where $\alpha_{em}$ is the fine-structure constant, $G_{F}$ is the Fermi coupling constant, $\left|V_{cd}\right|$ is the Cabibbo-Kobayashi-Maskawa matrix element [11, 12], $m_{D^{+}}$ is the nominal mass of the $D^{+}$ meson [13], and $F_{V}$ and $F_{A}$ are the form factors characterizing the non-perturbative quantum chromodynamics (QCD) dynamics. Various theoretical models have been proposed to investigate the $D^{+}\to\gamma e^{+}\nu_{e}$ decay, including quark models [14, 15, 16], perturbative QCD [17], lattice QCD [Desiderio:2020oej], and QCD factorization approaches [18, 19, 10]. Most theoretical predictions on its BF, $\mathcal{B}(D^{+}\to\gamma e^{+}\nu_{e})$, are on the order of $10^{-5}$, as summarized in Table 1.
@@
Theoretical predictions and experimental result for $\mathcal{B}(D^{+}\to\gamma e^{+}\nu_{e})$. Model $\mathcal{B}\ (\times 10^{-5})$ Light front quark model [14] 0.69 Non-relativistic quark model [15] 0.46 Relativistic quark model [16] 3.34 Perturbative QCD [17] $8.2\pm 6.5$ Lattice QCD [Desiderio:2020oej] $0.09\pm 0.04$ QCD factorization [18] 2.81 QCD factorization [19] 1.92 QCD factorization [10] $\left(1.88^{+0.36}_{-0.29},\,2.31^{+0.65}_{-0.54}\right)$ BESIII 2017 [20] <3.0<3.0
The only experimental result of $D^{+}\to\gamma e^{+}\nu_{e}$ was reported from the BESIII experiment [20], based on 2.9 $\rm fb^{-1}$ of $e^{+}e^{-}$ annihilation data taken at the center-of-mass energy $\sqrt{s}=3.773\penalty 10000\ \,\mathrm{GeV}$. The upper limit on the partial BF of $D^{+}\to\gamma e^{+}\nu_{e}$ with $E_{\gamma}>10$ MeV was determined to be $3.0\times 10^{-5}$ at 90% confidence level (C.L.), approaching the range of theoretical predictions [16, 17, 18, 19, 10]. A larger dataset comprising 20.3 $\rm fb^{-1}$ of $e^{+}e^{-}$ annihilation recently collected at BESIII [21, 22, 23], which includes the previous 2.9 $\rm fb^{-1}$, provides an excellent opportunity to further constrain these theoretical models.
@@
An unprecedentedly powerful signal identification tool can be accomplished using deep learning, utilizing a deep neural network (DNN), to distinguish signals from backgrounds. Leveraging flexible data representation and modern algorithms, this method exhibits powerful capabilities in recognizing and interpreting underlying relationships and hidden patterns within the topological structure of $e^{+}e^{-}$ annihilation events.
In this paper, we present an improved search for $D^{+}\to\gamma e^{+}\nu_{e}$ based on the 20.3 $\mathrm{fb}^{-1}$ data, enhancing signal sensitivity through a deep-learning-based signal identification method. Throughout this paper, charge conjugation is implied by default.
@@ 2 BESIII detector and Monte Carlo simulation
The BESIII detector [24] records symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring [25] in the center-of-mass energy range from 1.84 to 4.95 GeV, with a peak luminosity of $1.1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $\sqrt{s}=3.773\;\text{GeV}$. BESIII has collected large data samples in this energy region [26, 27, 28]. The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identification modules interleaved with steel. The charged-particle momentum resolution at $1\penalty 10000\ {\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the plastic scintillator TOF barrel region is 68 ps, while that in the end cap region was 110 ps. The end cap TOF system was upgraded in 2015 using multigap resistive plate chamber technology, providing a time resolution of 60 ps, which benefits 85% of the data used in this analysis [29, 30, 31].
Monte Carlo (MC) simulated data samples produced with a geant4-based [32] software package, which includes the geometric description of the BESIII detector and the detector response, were used in this study to determine detection efficiencies and estimate backgrounds. A signal MC sample of the decay $D^{+}\to\gamma e^{+}\nu_{e}$ was simulated according to Eq. (1), where the form factors were taken from Ref. [19]. The minimum energy of the radiative photon was set at 10 MeV to avoid the infrared divergence for soft photons. In the simulation, we modeled the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc [33, 34]. The inclusive MC sample encompassed the production of $D\bar{D}$ pairs (including quantum coherence for the neutral $D$ channels), non-$D\bar{D}$ decays of the $\psi(3770)$, ISR production of the $J/\psi$ and $\psi(3686)$ states, and continuum processes incorporated in kkmc [33, 34]. All particle decays were modeled with evtgen [35, 36] using BFs either taken from the Particle Data Group [13], when available, or otherwise estimated with lundcharm [37, 38]. Final state radiation from charged final state particles was incorporated using the photos package [39].
@@ 3 Initial data analysis
Taking advantage of the $D^{+}D^{-}$ pair production in $e^{+}e^{-}$ annihilation at $\sqrt{s}=3.773$ GeV, we employed a double-tag (DT) analysis method [40] to constrain the kinematics of $D^{+}\to\gamma e^{+}\nu_{e}$ and suppress non-$D^{+}D^{-}$ backgrounds. The $D^{-}$ meson was firstly reconstructed through six specific hadronic decay modes, $K^{+}\pi^{-}\pi^{-}$, $K^{+}\pi^{-}\pi^{-}\pi^{0}$, $K^{0}_{S}\pi^{-}$, $K^{0}_{S}\pi^{-}\pi^{0}$, $K^{0}_{S}\pi^{+}\pi^{-}\pi^{-}$ and $K^{+}K^{-}\pi^{-}$. Here, $K^{0}_{S}$ was reconstructed with a $\pi^{+}\pi^{-}$ pair, and $\pi^{0}$ was reconstructed from two photons. These decay modes are referred to as tag modes, and the selected events are called single-tag (ST) events. In the presence of ST $D^{-}$ meson, the signal decay $D^{+}\to\gamma e^{+}\nu_{e}$ was searched for within its recoil system, with selected events referred to as DT events. The BF of the signal decay was determined from
@@
where $N_{\rm DT}$ represents the sum of DT yields for all tag modes, and $N^{i}_{\rm ST}$, $\epsilon^{i}_{\rm ST}$, and $\epsilon^{i}_{\rm DT}$ are the ST yield, ST efficiency, and DT efficiency for the tag mode $i$, respectively.
For reconstructing the ST $D^{-}$ meson, charged tracks detected in the MDC were required to be within a polar angle ($\theta$) range of $|\rm{cos\theta}|<0.93$, where $\theta$ is defined with respect to the $z$-axis, which is the symmetry axis of the MDC. For charged tracks not originating from $K_{S}^{0}$ decays, the distance of closest approach to the interaction point (IP) was required to be less than 10 cm along the $z$-axis, $V_{z}$, and less than 1 cm in the transverse plane, $V_{xy}$. Particle identification (PID) for charged tracks combines measurements of the energy loss in the MDC (dEE/dxx) and the flight time in the TOF to form likelihoods $\mathcal{L}(h)\penalty 10000\ (h=K,\pi)$ for each hadron $h$ hypothesis. The charged kaons or pions not originating from $K_{S}^{0}$ were required to satisfy $\mathcal{L}(K)>\mathcal{L}(\pi)$ or $\mathcal{L}(\pi)>\mathcal{L}(K)$, respectively.
The $K_{S}^{0}$ candidates were reconstructed from two oppositely charged tracks satisfying $V_{z}<$ 20 cm. The two charged tracks were assigned as $\pi^{+}\pi^{-}$ without imposing further PID criteria. They were constrained to originate from a common vertex and required to have an invariant mass in the range of (487, 511) $\mathrm{MeV}/c^{2}$ corresponding to approximately $\pm 3\sigma$ of the detection resolution. The decay length of the $K^{0}_{S}$ candidate was required to be greater than twice the vertex resolution away from the IP.
Photon candidates were identified using showers in the EMC. The deposited energy of each shower was required to exceed 25 MeV in the barrel region ($|\cos\theta|<0.80$) and 50 MeV in the end cap region ($0.86<|\cos\theta|<0.92$). To exclude showers that originated from charged tracks, the angle subtended by the EMC shower and the position of the closest charged track at the EMC was required to exceed 10 degrees as measured from the IP. To suppress electronic noise and showers unrelated to the event, the difference between the EMC time and the event start time was required to be within [0, 700] ns. The $\pi^{0}$ mesons were reconstructed using a kinematic fit from all combinations of two photon candidates by constraining their invariant mass to the nominal $\pi^{0}$ mass [13]. The invariant mass of two photons before the kinematic fit was required to be in the range of (115, 150) $\mathrm{MeV}/c^{2}$. The combinations with both photons from end cap EMC regions are rejected because of poor resolution.
After selecting the component particles, we reconstruct the ST $D^{-}$ meson using their combinations. Two kinematic variables, the energy difference $\Delta E$, and the beam-constrained mass $M_{\rm BC}$ were calculated for signal extraction as follows:
@@
where $E_{\rm beam}$ is the beam energy, and $\vec{p}_{D^{-}}$ and $E_{D^{-}}$ are the momentum vector and energy of $D^{-}$ candidate in the $e^{+}e^{-}$ rest frame, respectively. If multiple combinations were present in an event, the combination with the smallest $|\Delta E|$ was retained for further analysis for each tag mode and charge. To suppress combinatorial backgrounds, mode-dependent requirements on $\Delta E$ were imposed on the accepted candidates within an approximate range of $\pm 3.5\sigma$ for the detection resolution, as listed in Table 3.
@@
$\Delta E$ requirements, ST yields in data, ST and DT efficiencies for each tag mode in data. Note the efficiencies do not include the BFs of subsequent $\pi^{0}$ and $K^{0}_{S}$ decays. Tag mode $\Delta E$ (MeV) $N_{\rm{ST}}^{i}\ (\times 10^{3})$ $\epsilon_{\rm{ST}}^{i}\ (\%)$ $\epsilon_{\rm{DT}}^{i}\ (\%)$ $K^{+}\pi^{-}\pi^{-}$ (25, 24)(-25,\ 24) $5552.8\pm 2.5$ $51.10$ 10.78 $K^{0}_{S}\pi^{-}$ (25, 26)(-25,\ 26) $656.5\pm 0.8$ $51.42$ 10.69 $K^{+}\pi^{-}\pi^{-}\pi^{0}$ (57, 46)(-57,\ 46) $1723.7\pm 1.8$ $24.40$ 5.42 $K^{0}_{S}\pi^{-}\pi^{0}$ (62, 49)(-62,\ 49) $1442.4\pm 1.5$ $26.45$ 5.57 $K^{0}_{S}\pi^{-}\pi^{-}\pi^{+}$ (28, 27)(-28,\ 27) $790.2\pm 1.1$ $29.59$ 6.62 $K^{+}K^{-}\pi^{-}$ (24, 23)(-24,\ 23) $481.4\pm 0.9$ $40.91$ 8.96
The ST yields were obtained through maximum likelihood fits to the $M_{\rm BC}$ distributions after applying the aforementioned requirements, as shown in Figure 3. In each fit, the signal was described by the shape obtained from MC simulation, convolved with a double-Gaussian function to take into account resolution differences between data and MC simulation results. The background was described by an ARGUS function [41] with the endpoint fixed at the beam energy and other parameters left free. A signal region $M_{\rm BC}\in(1.863,\ 1.877)\ \,\mathrm{GeV}/c^{2}$, corresponding to approximately $\pm 3\sigma$ of the detection resolution, was set to further reject the backgrounds. The ST efficiencies were estimated by analyzing the inclusive MC sample with the same method as in data, dividing the fitted yields by the numbers of generated events. Table 3 summarizes the ST yields in data and ST efficiencies for each tag mode.
@@
The signal candidates for $D^{+}\to\gamma e^{+}\nu_{e}$ were selected from the remaining charged tracks and showers, excluding those associated with the ST $D^{-}$ meson. We required only one track left with charge opposite to that of the ST $D^{-}$ meson, following the same selection criteria as those in the tag side. This track was required to be identified as a positron combining the PID information from $\mathrm{d}E/\mathrm{d}x$, the TOF system, and the EMC, satisfying $\mathcal{L}(e)>0.001$ and $\mathcal{L}(e)/[\mathcal{L}(e)+\mathcal{L}(\pi)+\mathcal{L}(K)]>0.8$ for the combined confidence levels. We also required that at least one photon candidate satisfied the same selection criteria used in the tag side. If multiple photon candidates passed the selection, the most energetic one was chosen as the signal radiative photon. To improve the kinematic resolution of the positron, final-state-radiation recovery was performed by adding the momenta of any remaining photon candidates back to the reconstructed momentum of positron candidate if their opening angles with respect to the IP were less than $5^{\circ}$.
To derive information about the missing neutrino, a kinematic observable $U_{\rm miss}$ was defined as
@@
where $E_{\rm miss}$ and $\vec{p}_{\rm miss}$ represent the total energy and momentum vector of all missing particles in the event. The missing energy was calculated as $E_{\rm miss}=E_{\rm beam}-E_{e^{+}}-E_{\gamma}$, where $E_{e^{+}}$ and $E_{\gamma}$ are the energies of signal positron and photon, respectively. The missing momentum was calculated as $\vec{p}_{\rm miss}=\vec{p}_{D^{+}}-\vec{p}_{e^{+}}-\vec{p}_{\gamma}$, where $\vec{p}_{D^{+}}$, $\vec{p}_{e^{+}}$ and $\vec{p}_{\gamma}$ are the momenta of signal $D^{+}$ meson, positron, and photon, respectively. The $D^{+}$ momentum was further constrained as $\vec{p}_{D^{+}}=-\frac{\vec{p}_{\rm ST}}{\left|\vec{p}_{\rm ST}\right|}\sqrt{E^{2}_{\rm beam}-m^{2}_{D^{+}}}$, where $\vec{p}_{\rm ST}$ is the momentum of ST $D^{-}$ meson. The consequent analysis was performed within $U_{\rm miss}\in(-0.2,\ 0.2)$ GeV.
We also explored $D^{+}\to\pi^{0}e^{+}\nu_{e}$ as a validation channel in this study. For reconstructing $D^{+}\to\pi^{0}e^{+}\nu_{e}$ events, additional DT selection criteria were applied after those mentioned above. At least one $\pi^{0}$ candidate was required, and the one with minimum $\chi^{2}$ from the kinematic fit was chosen as the signal $\pi^{0}$ candidate. Another kinematic observable, $U^{\prime}_{\rm miss}$, was defined as
where $E^{\prime}_{\rm miss}=E_{\rm beam}-E_{e^{+}}-E_{\pi^{0}}$ and $\vec{p^{\prime}}_{\rm miss}=\vec{p}_{D^{+}}-\vec{p}_{e^{+}}-\vec{p}_{\pi^{0}}$ were calculated using the energy and momentum of the signal $\pi^{0}$ candidate, respectively.
Figure 3(a) presents the $U_{\rm miss}$ distributions for both data and inclusive MC samples after applying the above selections. The major backgrounds are from $D^{+}\to\pi^{0}e^{+}\nu_{e}$, $D^{+}\to K^{0}_{L}e^{+}\nu_{e}$, and $D^{+}\to K^{0}_{S}(\to\pi^{0}\pi^{0})e^{+}\nu_{e}$. Notably, the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background
@@
(a) $U_{\rm miss}$ distribution of the signal decay $D^{+}\to\gamma e^{+}\nu_{e}$ and (b) $U^{\prime}_{\rm miss}$ distribution of the validation channel $D^{+}\to\pi^{0}e^{+}\nu_{e}$ after the DT selection. The black dots with error bars represent data. The histograms with filled colors represent different background components in the inclusive MC sample. The red hollow histogram indicates an arbitrarily scaled signal shape.
exhibits a non-trivial shape within the signal region. The high background level makes it difficult to reliably extract the signal yield from the $U_{\rm miss}$ distribution with adequate sensitivity, thereby requiring further signal identification treatments, as explained below.
4 Deep-learning-based signal identification
The main challenge in distinguishing the signals from massive backgrounds lies in extracting distinctive features from the data. Notably, the shower patterns observed in the EMC provide key information for this task. Background events involving a $\pi^{0}$ typically result in a higher number of photon candidates than the signal, whereas $K^{0}_{L}$-induced showers often exhibit more dispersed cluster shapes compared to those from photons [42]. Additionally, the kinematics of the particles, influenced by the unique phase spaces and dynamics of their decays, offer another aspect of differentiation. Conventional cut-based methods characterize such distinctions via hand-crafted variables derived from practical understanding of the involved physics, such as the number of $\pi^{0}$-like photon combinations, lateral moment of photon candidates [20], and $\chi^{2}$ values from kinematic fits under specific final-state hypotheses [43]. In contrast, deep learning methods aim to process minimally refined data at the level of fundamental detector responses to the penetrations of the final state particles. A DNN is expected to automatically explore underlying correlations and recognize distinctive information beyond prior knowledge.
In this study, the dataset for DNN training was constructed by randomly shuffling signal and background events sourced from relevant MC samples. The events were categorized into three groups with equal statistics: $D^{+}\to\gamma e^{+}\nu_{e}$ signal, $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background, and other inclusive backgrounds. The training dataset contained approximate 3.3 million events in total, with 80% used for training and the remaining 20% reserved for validation. The DNN input information from an event included all charged tracks reconstructed in the MDC and isolated showers clustered in the EMC. These particles were organized as a point cloud [44] structure, i.e., an unordered and variable-sized set of points in a high-dimensional feature space. For each charged track, the features comprised the azimuthal and polar angles in the center-of-mass frame, charge, magnitude of momentum, and parameters characterizing its helical trajectory in the MDC. Additionally, low-level measurements from the MDC, TOF, and EMC were incorporated as implicit information for particle identification and reconstruction quality. For each shower, the features comprised the azimuthal and polar angles, energy deposition, count of fired crystals, time measurement, and parameters describing its expansion scope among nearby crystals. A full list of input features is provided in Table 4; these features were found to be consistent between data and MC simulation results.